If three coins are tossed. Find the probability that no more than one coin lands heads up. Is it 5/8. 1/4, 3/8, or 1/2
Please help.

Answer by stanbon(48568)

If three coins are tossed. Find the probability that no more than one coin lands heads up. Is it 5/8. 1/4, 3/8, or 1/2
--------------------
P(0 or 1 heads in 3 tosses) = P(no heads) + P(1 head)
= (1/2)^3 + 3C1(1/2)^1*(1/2)^2
= (1/8) + 3*(1/8) = 4/8 = 1/2
=================================